OpenStudy (anonymous):

Find the derivative of the function: f(x)=4arcsin(x/2)

OpenStudy (anonymous):

I did some work and got: \[\frac{ 4\sqrt{4-x^2} }{ 4-x^2 }\]

zepdrix (zepdrix):

\[\Large f(x)\quad=\quad 4\arcsin\left(\frac{x}{2}\right)\] \[\Large f'(x)\quad=\quad 4\left(\frac{1}{\sqrt{1-\left(\frac{x}{2}\right)^2}}\right)\left(\frac{1}{2}\right)\]

zepdrix (zepdrix):

And I guess you did some shinanigans to simplify it down? Hmm ok ok let's see..

zepdrix (zepdrix):

\[\Large f'(x)\quad=\quad 2\left(\frac{1}{\sqrt{1-\frac{x^2}{4}}}\right)\] Multiplying through by 2/2 gives us:\[\Large f'(x)\quad=\quad \frac{4}{\sqrt{4-x^2}}\]

zepdrix (zepdrix):

And then I guess you rationalized it, giving usssssss... Yah looks like you did it correctly! :)

OpenStudy (anonymous):

alright perfect. now if you dont mind helping me with one last question: find the derivative of g(x)=e^2xarctan(2x) I got up to: g'(x)=e^2x\[\frac{ 2 }{ 1+4x^2 }\]

OpenStudy (anonymous):

g'(x)=e^2x*2/1+4x^2 I dont know what to do from here

zepdrix (zepdrix):

Looks like we'll need to start with product rule, yes? :)\[\Large g(x)=e^2x \arctan(2x)\] \[\Large g'(x)=\color{royalblue}{(e^{2x})'}\arctan2x+e^{2x}\color{royalblue}{(\arctan2x)'}\]

zepdrix (zepdrix):

The blue parts are where we need to take a derivative. Confusion? :u

OpenStudy (anonymous):

I understand. This is what I get: e^2x(2) * [2/(1+2x)]+e^2x[4/(1+2x)^2]

OpenStudy (anonymous):

or is it: (2) * e^2x * [2/1+4x^2]+ e^2x [16x/(1+4x^2)^2]

zepdrix (zepdrix):

ah sorry i ran off to make some food :d mmm let's see...

zepdrix (zepdrix):

\[\Large g'(x)\quad =\quad\color{royalblue}{(e^{2x})'}\arctan2x\quad+\quad e^{2x}\color{royalblue}{(\arctan2x)'}\] \[\Large g'(x)\quad=\quad\color{orangered}{2e^{2x}}\arctan2x\quad+\quad e^{2x}\color{orangered}{\frac{2}{1+(2x)^2}}\]

zepdrix (zepdrix):

We should only be differentiating the arctan within the second term, right? :)

zepdrix (zepdrix):

That arctan derivative giving you a little trouble?